OpenStudy (anonymous):

when i did 2/3+5/6 my anwser was 7/13 help again

OpenStudy (anonymous):

any1 to anwser?

OpenStudy (anonymous):

It should be 3/2

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

so much

OpenStudy (anonymous):

Try this method: find the LCD of both denominators which is 18 and cross multiply both fractions and add them you'll get 20/18 which is 3/2

OpenStudy (lalaly):

\[(\frac{2}{3} \times \frac{2}{2}) + \frac{5}{6} = \frac{4}{6}+\frac{5}{6} =\frac{9}{6}=\frac{3}{2}\]

OpenStudy (anonymous):

does any1 know what 3/4 +2/3 is ?